# Gorenstein singularities with $\mathbb{G}_{m}$-action and moduli spaces of holomorphic differentials

### v1

@@ 1.6. Affine and tautological geometry of subcanonical points
Finally, we turn to the affine and tautological geometry of moduli spaces of differentials. It was shown in [Che19a] that the strata of (strictly) meromorphic canonical divisors with prescribed zero and pole orders are affine varieties by using relations of tautological divisor classes in the strata. However, the same question remains open for the holomorphic case. Therefore, it is meaningful to study the affine and tautological geometry for each $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ in the decomposition of the strata $\mathbb{P}\mathcal{H}(\mu)$ of holomorphic differentials, and the results can help us better understand the global geometry of $\mathbb{P}\mathcal{H}(\mu)$. Here we focus on the case of $(Y,q)$ being a monomial singularity, where $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ can be identified with the locus $\mathcal{M}_{g,1}^{H}$ of Weierstrass points $(C,p)$ with a fixed semigroup $H$. In particular, $2g-1\not\in H$ is equivalent to that $(2g-2)p$ is a canonical divisor, i.e., $(C,p)\in\mathbb{P}\mathcal{H}(2g-2)$. Such a semigroup $H$ is called symmetric and such a point $p$ is called a subcanonical point.
@@
In Section 8, we provide more detailed descriptions for the above results in Theorem 8.3 and Theorem 8.5. Additionally, we apply the results to study the varying minimal spin strata in genus six; see Example 8.7 and Example 8.8.
@@ 1.7. Further directions
The geometry of $\mathbb{P}\mathcal{H}(\mu)$ and its decomposition into deformation spaces $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ of Gorenstein singularities with $\mathbb{G}_{m}$-action can motivate a number of intriguing $K(\pi,1)$ and affine patching related questions; see [Kon97, LM14, Mon17, Gia24] for some related conjectures and results. In particular, our work has identified each of the nonvarying strata with the locus of smooth deformations of the corresponding singularity, which can be regarded as the discriminant hypersurface complement in the versal deformation space. This viewpoint has played a crucial role in the application of the $K(\pi,1)$-property for $ADE$ singularities established by Brieskorn and Deligne; see [Bri73, Del72] and also [Loo84a, Chapter 9] for a discussion whose viewpoint is closer to ours. Therefore, the explicit defining equations of those nonvarying singularities we have obtained pave an avenue towards determining the $K(\pi,1)$-property for the corresponding strata of differentials.
Additionally, certain nonreduced Gorenstein curves (such as ribbons) can naturally appear in the log minimal model program for ¯$\overline{\mathcal{M}}_{g}$; see [AFS16, Section 3.3]. Therefore, it would be interesting to find a correspondence between holomorphic differentials and nonreduced Gorenstein curves. Similarly, we can ask how other singularities, e.g., those that are not Gorenstein, or without $\mathbb{G}_{m}$-action, or of higher dimension, can appear in the story, as well as how their deformation spaces fit together. From the perspective of differentials, we can also study other variants of the construction by using meromorphic differentials or higher-order differentials instead of holomorphic one-forms.
@@ 1.8. Outline of the paper
In Section 5, we study the singularities associated to the nonvarying strata of holomorphic differentials and prove Theorem 1.2. In the course of the proof, we also describe the invariants and defining equations for these singularities.
@@
In Section 7, we define the slopes of these singularities and utilize them from the viewpoint of bounding the slopes of effective divisors in ¯$\overline{\mathcal{M}}_{g}$, as in Theorem 1.5.
@@ Acknowledgements
The authors would like to thank Dan Abramovich, Luca Battistella, Sebastian Bozlee, Qile Chen, Yifei Chen, David Eisenbud, Yu-Wei Fan, Maksym Fedorchuk, Riccardo Giannini, Samuel Grushevsky, Brendan Hassett, David Holmes, Dihua Jiang, Hannah Larson, Yuchen Liu, Eduard Looijenga, Martin Möller, Adrian Neff, Dhruv Ranganathan, Song Sun, Jonathan Wise, Dan Xie, Junyan Zhao, Anton Zorich, Huaiqing Zuo, and Kang Zuo for inspiring discussions on many related topics.
@@ 2.1. Deformations of singularities with $\mathbb{G}_{m}$-action
Throughout the paper, we use $(Y,q)$ to denote (the analytic germ of) a reduced affine curve $Y$ with a unique isolated singularity at $q$. Suppose $(Y,q)$ admits a (good) $\mathbb{G}_{m}$-action that acts freely on $Y\setminus q$, with the convention that it acts with positive weights on the maximal ideal of $\widehat{\mathcal{O}}_{Y,q}$. In other words, $\widehat{\mathcal{O}}_{Y,q}\cong\mathbb{C}\{y_{1},\ldots,y_{n}\}/I$, where the weights $a_{i}$ of $y_{i}$ are either all positive or all negative and $I$ is a graded ideal generated by weighted homogeneous polynomials. The (good) $\mathbb{G}_{m}$-action on $(Y,q)$ is thus given by $\lambda\cdot(y_{1},\ldots,y_{n})=(\lambda^{a_{1}}y_{1},\ldots,\lambda^{a_{n}}y_{n})$. Such singularities are also called quasi-homogeneous.
Pinkham [Pin74, Section 2] showed that $(Y,q)$ has a $\mathbb{G}_{m}$-equivariant versal deformation space, which we denote by ${\rm Def}(Y,q)$. We remark that Pinkham’s definition of versal includes the extra condition that the tangent space to the base space of the deformation has minimal dimension, which is sometimes also called miniversal or semi-universal in the literature; see [Pin74, Definition (1.4)] and [Pin78, Section 2]. Here we follow Pinkham’s convention.
@@
has a natural grading. We denote by ${\rm Def}^{-}(Y,q)$ the subspace of ${\rm Def}(Y,q)$ corresponding to negatively graded deformations, i.e., the tangent space to ${\rm Def}^{-}(Y,q)$ is $T^{1,-}=\sum_{\nu<0}T^{1}(\nu)$. Both the family of negative deformations and the base space ${\rm Def}^{-}(Y,q)$ are defined by polynomials. By Pinkham’s convention, the maximal ideal of 00 in the coordinate ring of ${\rm Def}^{-}(Y,q)$ is positively graded due to the duality between the tangent vectors of $T^{1}$ and the deformation parameters of the base; see [Pin74, (2.9)]. We denote by
@@
In particular, if $\dim T^{1,-}=\dim{\rm Def}^{-}_{s}(Y,q)$ and $T^{2}=\{0\}$, then the weighted projective space $[({\rm Def}^{-}(Y,q)\setminus 0)/\mathbb{G}_{m}]\cong\mathbb{P}(T^{1,-})$ provides a compactification for the subset
$\mathbb{P}{\rm Def}^{-}_{s}(Y,q)\coloneqq[({\rm Def}^{-}_{s}(Y,q)\setminus 0)/\mathbb{G}_{m}]$ (1)
@@
For the reader’s convenience, we illustrate the above terms through the well-known example of the ordinary cusp $(Y,q)$ defined by $X_{1}^{3}-X_{2}^{2}$; see [Pin74, (1.16)]. The $\mathbb{G}_{m}$-action acts with weight $2$ on $X_{1}$ and weight $3$ on $X_{2}$. The versal deformation of the cusp is given by
@@
over $R=\mathbb{C}[t_{1},t_{2}]$. We assign to $t_{1}$ and $t_{2}$ the weights $4$ and $6$, respectively, and then $F$ becomes a homogeneous equation. In this case, $t_{1}$ and $t_{2}$ have positive weights, and hence by Pinkham’s convention, the versal deformation of the cusp has negative grading only. Next, we replace $t_{1}$ and $t_{2}$ by $t_{1}X_{0}^{4}$ and $t_{2}X_{0}^{6}$, respectively, where $X_{0}$ is a new variable with weight $1$. We obtain
@@
where $R[X_{0},X_{1},X_{2}]/(\overline{F})$ is graded in the $X_{i}$’s only and $\mathbb{G}_{m}$ acts on $R$ as before. It is easy to see that all fibers above points of the same $\mathbb{G}_{m}$-orbit in ${\rm Def}^{-}(Y,q)$ are isomorphic. Indeed, all fibers are smooth elliptic curves, except the original cuspidal curve $(Y,q)$ above $(t_{1},t_{2})$ and a family of rational nodal curves above a unique $\mathbb{G}_{m}$-orbit in ${\rm Def}^{-}(Y,q)$. Additionally, there is a natural section obtained by setting $X_{0}=0$. Therefore, the quotient $[({\rm Def}^{-}(Y,q)\setminus 0)/\mathbb{G}_{m}]$ thus provides the usual compactification of the affine $j$-line for parameterizing elliptic curves, where the boundary of the compactification parameterizes the isomorphism class of rational nodal curves.
@@
This example was already noted in [Pin74, (13.1)]. Consider the plane curve singularity $(Y,q)$ defined by $X_{1}^{5}-X_{2}^{4}$. It is a special case of monomial curve singularities whose associated semigroup is generated by $4$ and $5$; see Section 3.1 later for a detailed discussion on monomial curve singularities. Consider the deformation $X_{1}^{5}-X_{2}^{4}+tX_{1}^{3}X_{2}^{2}$. In this case, $X_{1}$ and $X_{2}$ have weights $4$ and $5$ under the $\mathbb{G}_{m}$-action, respectively. Therefore, the deformation parameter $t$ has weight $-2$. In other words, the corresponding tangent vector in $T^{1}$ has positive grading $2$, hence not in ${\rm Def}^{-}(Y,q)$. As a result, the $\mathbb{G}_{m}$-action along this direction is repulsive, which is not a good $\mathbb{G}_{m}$-action.
@@
$Y_{\infty}\coloneqq\overline{Y}-Y,$ (2)
@@
Looijenga showed that the part of the negatively graded versal deformation ${\rm Def}^{-}(Y,q)$ is a final object in the category of deformations of $(Y,q)$ with good $\mathbb{G}_{m}$-action; see [Loo84b, (A.2) and (A.6)]. Furthermore, for $X\in\mathbb{P}{\rm Def}^{-}(Y,q)$ with the divisor at infinity $X_{\infty}$, we have the isomorphism of graded algebras $R_{\overline{X}}/tR_{\overline{X}}\cong R$, where $R_{\overline{X}}\cong\bigoplus_{k=0}^{\infty}H^{0}(\mathcal{O}_{\overline{X}}(kX_{\infty}))$ and $t\in H^{0}(\mathcal{O}_{\overline{X}}(X_{\infty}))$ is the element corresponding to $1$; see [Loo84b, (A.5)].
@@
In general, ${\rm Def}(Y,q)$ can be reducible. An irreducible component of ${\rm Def}(Y,q)$ is called a smoothing component if its generic fiber is smooth. If $(Y,q)$ is a smoothable reduced curve singularity, then all smoothing components of $(Y,q)$ have the same dimension, which is called the Deligne number. Moreover, if $(Y,q)$ is an isolated Gorenstein curve singularity with good $\mathbb{G}_{m}$-action (i.e., quasi-homogeneous), then the Deligne number is equal to the Milnor number $12g+n-$ 1, where $g$ is the arithmetic genus of the projectivization of $Y$ and $n$ is the number of branches. We refer to [Gre20, Section 2.6] for a survey about these concepts and their relations.
@@ 2.2. From differentials to singularities
Let $\mu=(m_{1},\ldots,m_{n})$ be a partition of $22g-$ 2, where $m_{i}\geq 0$ for all $i$. Given any $(C,p_{1},\ldots,p_{n})\in\mathbb{P}\mathcal{H}(\mu)$, we will associate to it an isolated curve singularity. To this end, we define the following parameters that will be used frequently throughout the paper:
@@
~$\tilde{\mathcal{L}}\coloneqq\omega_{\tilde{\pi}}\left(\sum_{i=1}^{n}\tilde{P}_{i}+\ell\tilde{C}_{0}\right).$ (3)
@@
In particular, this family $\tilde{\pi}$ parameterizes an isotrivial degeneration of holomorphic differentials of type $\mu$ on $C$ to a multi-scale differential $\tilde{\eta}$ on $\tilde{C}_{0}$ in the sense of [BCG+18, BCG+24, CGH+22], where $\tilde{\eta}$ has a pole of order $m_{i}+2$ at the nodal branch along each $E_{i}$. By using the explanation in [CC24, Section 2], we thus obtain a birational contraction
$f\colon\tilde{\mathcal{C}}\longrightarrow\mathcal{C}\coloneqq\operatorname{Proj}\left(\bigoplus_{k}\tilde{\pi}_{*}\tilde{\mathcal{L}}^{k}\right)$ (4)
such that $f$ is an isomorphism away from $\tilde{C}_{0}$ and the images of $E_{1},\ldots,E_{n}$ intersect at a singularity formed by the contraction image of $\tilde{C}_{0}$. Additionally, $\tilde{\mathcal{L}}$ descends to a (relatively) ample line bundle $\mathcal{L}$ on $\mathcal{C}$, where $f^{*}\mathcal{L}\cong\tilde{\mathcal{L}}$.
@@
Later in Theorem 2.6 (iv), we will see that $\mathbb{P}\mathcal{H}(\mu)_{(Y,q)}$ can be identified with the locus of negatively graded smooth deformations $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ in the versal deformation space as introduced in (1).

### v2

@@ 1.7. Affine and tautological geometry of subcanonical points
Finally, we turn to the affine and tautological geometry of moduli spaces of differentials. It was shown in [Che19a] that the strata of (strictly) meromorphic canonical divisors with prescribed zero and pole orders are affine varieties by using relations of tautological divisor classes in the strata. However, the same question remains open for the holomorphic case. Therefore, it is meaningful to study the affine and tautological geometry for each $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ in the decomposition of the strata $\mathbb{P}\mathcal{H}(\mu)$ of holomorphic differentials, and the results can help us better understand the global geometry of $\mathbb{P}\mathcal{H}(\mu)$.
Here we concentrate on the case of $(Y,q)$ being a monomial singularity, where $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ can be identified with the locus $\mathcal{M}_{g,1}^{H}$ of Weierstrass points $(C,p)$ with a fixed semigroup $H$. In particular, $2g-1\not\in H$ is equivalent to that (2g2)p(2g-2)p is a canonical divisor, i.e., $(C,p)\in\mathbb{P}\mathcal{H}(2g-2)$. Such a semigroup $H$ is called symmetric and such a point $p$ is called a subcanonical point.
@@
In Section 8, we provide more detailed descriptions for the above results in Theorem 8.3 and Theorem 8.5. Additionally, we apply the results to study the varying minimal spin strata in genus six; see Example 8.7 and Example $8.$
@@ 1.8. Further directions
The explicit defining equations of the nonvarying singularities we have obtained pave an avenue towards determining the $K(\pi,1)$-property for the corresponding strata of differentials as well as understanding special loci in Teichmüller dynamics. Note that the Coxeter–Dynkin diagrams of $ADE$ singularities and the corresponding curve systems also appear in the construction of Teichmüller curves; see [McM23, Section 6] for more details.
Additionally, certain nonreduced Gorenstein curves (such as ribbons) can naturally appear in the log minimal model program for $\overline{\mathcal{M}}_{g}$; see [AFS16, Section 3.3]. Therefore, it would be interesting to find a correspondence between holomorphic differentials and nonreduced Gorenstein curves. Similarly, we can ask how other singularities, e.g., those that are not Gorenstein, or without $\mathbb{G}_{m}$-action, or of higher dimension, can appear in the story, as well as how their deformation spaces fit together. From the perspective of differentials, we can also study other variants of the construction by using meromorphic differentials or higher-order differentials instead of holomorphic one-forms.
@@ 1.9. Outline of the paper
In Section 5, we study the singularities associated to the nonvarying strata of holomorphic differentials and prove Theorem 1.2. We also study the behavior of the $K(\pi,1)$-property for singularities obtained by adding an ordinary marked point, which proves Theorem 1.3. In the course of the proofs, we describe the invariants and defining equations for these singularities. Additionally, we study several varying strata and describe the interesting geometry for the singularities associated to them.
@@
In Section 7, we define the slopes of singularities and utilize them from the viewpoint of bounding the slopes of effective divisors in $\overline{\mathcal{M}}_{g}$, as in Theorem 1.6.
@@ Acknowledgements
The authors would like to thank Dan Abramovich, Jarod Alper, Luca Battistella, Sebastian Bozlee, Iacopo Brivio, Qile Chen, Yifei Chen, Giulio Codogni, David Eisenbud, Yu-Wei Fan, Gavril Farkas, Maksym Fedorchuk, Riccardo Giannini, Samuel Grushevsky, Paul Hacking, Brendan Hassett, David Holmes, Dihua Jiang, Bochao Kong, Hannah Larson, Yongnam Lee, Yuchen Liu, Eduard Looijenga, Michael Lönne, Martin Möller, Adrian Neff, Yuji Odaka, Yu Qiu, Dhruv Ranganathan, Yongbin Ruan, Song Sun, Luca Tasin, Filippo Viviani, Jonathan Wise, Alex Wright, Dan Xie, Junyan Zhao, Anton Zorich, Huaiqing Zuo, and Kang Zuo for inspiring discussions on many related topics.
@@ 2.1. Deformations of singularities with $\mathbb{G}_{m}$-action
Throughout the paper, we use $(Y,q)$ to denote (the analytic germ of) a reduced affine curve $Y$ with a unique isolated singularity at $q$. Suppose $(Y,q)$ admits a (good) $\mathbb{G}_{m}$-action that acts freely on $Y\setminus q$, with the convention that it acts with all positive or all negative weights on the maximal ideal of $\widehat{\mathcal{O}}_{Y,q}$. In other words, $\widehat{\mathcal{O}}_{Y,q}\cong\mathbb{C}\{y_{1},\ldots,y_{n}\}/I$, where the weights $a_{i}$ of $y_{i}$ are either all positive or all negative and $I$ is a graded ideal generated by weighted homogeneous polynomials. The (good) $\mathbb{G}_{m}$-action on $(Y,q)$ is thus given by $\lambda\cdot(y_{1},\ldots,y_{n})=(\lambda^{a_{1}}y_{1},\ldots,\lambda^{a_{n}}y_{n})$. Such singularities are also called quasi-homogeneous.
Pinkham [Pin74, Section 2] showed that $(Y,q)$ has a $\mathbb{G}_{m}$-equivariant versal deformation space, which we denote by ${\rm Def}(Y,q)$. We remark that Pinkham’s definition of “versal” includes the extra condition that the tangent space to the base space of the deformation has minimal dimension, which is commonly called miniversal or semi-universal in the literature; see [Pin74, Definition (1.4)] and [Pin78, Section 2].
@@
has a natural grading. We denote by ${\rm Def}^{-}(Y,q)$ the subspace of ${\rm Def}(Y,q)$ corresponding to negatively graded deformations, i.e., the tangent space to ${\rm Def}^{-}(Y,q)$ is $T^{1,-}=\sum_{\nu<0}T^{1}(\nu)$. Both the family of negative deformations and the base space ${\rm Def}^{-}(Y,q)$ are defined by polynomials. By Pinkham’s convention, the maximal ideal of 0 in the coordinate ring of ${\rm Def}^{-}(Y,q)$ is positively graded due to the duality between the tangent vectors of $T^{1}$ and the deformation parameters of the base; see [Pin74, (2.9)]. We denote by
@@
In particular, if $\dim T^{1,-}=\dim{\rm Def}^{-}_{s}(Y,q)$ and $T^{2}=\{0\}$, then the weighted projective space $({\rm Def}^{-}(Y,q)\setminus 0)/\mathbb{G}_{m}]\cong\mathbb{P}(T^{1,-})$ provides a compactification for the subset
(1) $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)\coloneqq[({\rm Def}^{-}_{s}(Y,q)\setminus 0)/\mathbb{G}_{m}]$
@@
For the reader’s convenience, we illustrate the above terms through the well-known example of the ordinary cusp $(Y,q)$ defined by $X_{1}^{3}-X_{2}^{2}$; see [Pin74, (1.16)]. The $\mathbb{G}_{m}$-action acts with weight $2$ on $X_{1}$ and weight $3$ on $X_{2}$. The miniversal deformation of the cusp is given by
@@
over $R=\mathbb{C}[t_{1},t_{2}]$. We assign to $t_{1}$ and $t_{2}$ the weights $4$ and $6$, respectively, and then $F$ becomes a homogeneous equation. In this case, $t_{1}$ and $t_{2}$ have positive weights, and hence by Pinkham’s convention, the miniversal deformation of the cusp has negative grading only. Next, we replace $t_{1}$ and $t_{2}$ by $t_{1}X_{0}^{4}$ and $t_{2}X_{0}^{6}$, respectively, where $X_{0}$ is a new variable with weight $1$. We obtain
@@
where $R[X_{0},X_{1},X_{2}]/(\overline{F})$ is graded in the $X_{i}$’s only and $\mathbb{G}_{m}$ acts on $R$ as before. It is easy to see that all fibers above points of the same $\mathbb{G}_{m}$-orbit in ${\rm Def}^{-}(Y,q)$ are isomorphic. Indeed, all fibers are smooth elliptic curves, except the original cuspidal curve $(Y,q)$ above $(t_{1},t_{2})$ and a family of rational nodal curves above a unique $\mathbb{G}_{m}$-orbit in ${\rm Def}^{-}(Y,q)$. Additionally, there is a natural section obtained by setting $X_{0}=0$. Therefore, the quotient $({\rm Def}^{-}(Y,q)\setminus 0)/\mathbb{G}_{m}]$ thus provides the usual compactification of the affine $j$-line for parameterizing elliptic curves, where the boundary of the compactification parameterizes the isomorphism class of rational nodal curves.
@@
This example was already noted in [Pin74, (13.1)]. Consider the plane curve singularity $(Y,q)$ defined by $X_{1}^{5}-X_{2}^{4}$. It is a special case of monomial curve singularities whose associated semigroup is generated by $4$ and $5$; see Section 3.1 later for a detailed discussion on monomial curve singularities. Consider the deformation $X_{1}^{5}-X_{2}^{4}+tX_{1}^{3}X_{2}^{2}$. In this case, $X_{1}$ and $X_{2}$ have weights $4$ and $5$ under the $\mathbb{G}_{m}$-action, respectively. Therefore, the deformation parameter $t$ has weight 2-2. In other words, the corresponding tangent vector in $T^{1}$ has positive grading $2$, hence not in ${\rm Def}^{-}(Y,q)$. As a result, the $\mathbb{G}_{m}$-action along this direction is repulsive, which is not a good $\mathbb{G}_{m}$-action.
@@
(2) $Y_{\infty}\coloneqq\overline{Y}-Y,$
@@
Looijenga showed that the part of the negatively graded miniversal deformation ${\rm Def}^{-}(Y,q)$ is a final object in the category of deformations of $(Y,q)$ with good $\mathbb{G}_{m}$-action; see [Loo84b, (A.2) and (A.6)]. Furthermore, for $X\in\mathbb{P}{\rm Def}^{-}(Y,q)$ with the divisor at infinity $X_{\infty}$, we have the isomorphism of graded algebras $R_{\overline{X}}/tR_{\overline{X}}\cong R$, where $R_{\overline{X}}\cong\bigoplus_{k=0}^{\infty}H^{0}(\mathcal{O}_{\overline{X}}(kX_{\infty}))$ and $t\in H^{0}(\mathcal{O}_{\overline{X}}(X_{\infty}))$ is the element corresponding to $1$; see [Loo84b, (A.5)].
@@
In general, ${\rm Def}(Y,q)$ can be reducible. An irreducible component of ${\rm Def}(Y,q)$ is called a smoothing component if its generic fiber is smooth. If $(Y,q)$ is a smoothable reduced curve singularity, then all smoothing components of $(Y,q)$ have the same dimension, which is called the Deligne number. Moreover, if $(Y,q)$ is an isolated Gorenstein curve singularity with good $\mathbb{G}_{m}$-action (i.e., quasi-homogeneous), then the Deligne number is equal to the Milnor number 2g+n12g+n-1, where $g$ is the arithmetic genus of the projectivization of $Y$ and $n$ is the number of branches. We refer to [Gre20, Section 2.6] for a survey about these concepts and their relations.
@@ 2.2. From differentials to singularities
Let $\mu=(m_{1},\ldots,m_{n})$ be a partition of 2g22g-2, where $m_{i}\geq 0$ for all $i$. Given any $(C,p_{1},\ldots,p_{n})\in\mathbb{P}\mathcal{H}(\mu)$, we will associate to it an isolated curve singularity. To this end, we define the following parameters that will be used frequently throughout the paper:
@@
(3) $\tilde{\mathcal{L}}\coloneqq\omega_{\tilde{\pi}}\left(\sum_{i=1}^{n}\tilde{P}_{i}+\ell\tilde{C}_{0}\right).$
@@
In particular, this family $\tilde{\pi}$ parameterizes an isotrivial degeneration of holomorphic differentials of type $\mu$ on $C$ to a multi-scale differential $\tilde{\eta}$ on $\tilde{C}_{0}$ in the sense of [BCG+18, BCG+24, CGH+25], where $\tilde{\eta}$ has a pole of order $m_{i}+2$ at the nodal branch along each $E_{i}$. By using the explanation in [CC26, Section 2], we can contract $\tilde{C}_{0}$ through the following birational contraction
(4) $f\colon\tilde{\mathcal{C}}\longrightarrow\mathcal{C}\coloneqq\operatorname{Proj}\left(\bigoplus_{k}\tilde{\pi}_{*}\tilde{\mathcal{L}}^{k}\right)$
where $f$ is an isomorphism away from $\tilde{C}_{0}$ and the images of $E_{1},\ldots,E_{n}$ intersect at a singularity formed by the contraction image of $\tilde{C}_{0}$; see also [Smy13, Proposition 2.6] for a more general discussion of contractions. Additionally, $\tilde{\mathcal{L}}$ descends to a (relatively) ample line bundle $\mathcal{L}$ on $\mathcal{C}$, where $f^{*}\mathcal{L}\cong\tilde{\mathcal{L}}$.
@@
Later in Theorem 2.6 (iv), we will see that $\mathbb{P}\mathcal{H}(\mu)_{(Y,q)}$ can be identified with the locus of negatively graded smooth deformations $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ in the miniversal deformation space as introduced in (1).